12. Given that 2.V 3 which of the following choices will make S linearly independent? b) v= 1-15 24 0 d) v

Please show me the detail using matrix form.
Answer #1

if the matrix formed by the vectors have determinant as 0, then vectors are linear dependent.

(a) A = 1-5-2-7 | , det(A) = 0 as C3 = C2+ C so vectors are linearly dependent

\\(b)\ A = \begin{bmatrix} 2 &1 &6 \\ -5 & -2 &-15 \\ 8& 3& 24 \end{bmatrix},\ det(A) =0 \\ \\so\ vectors\ are\ linearly\ dependent\\

-5-2-151-2x-2-15 8 3 24 15 +5/ 24 =2x13 24 2 -15

\\(c)\ A = \begin{bmatrix} 2 &1 &2 \\ -5 & -2 &-1 \\ 8& 3& 5 \end{bmatrix},\ det(A) =5\\ \\so\ vectors\ are\ linearly\ independent\\

1 2 +8x 2 -1

\\(d)\ A = \begin{bmatrix} 2 &1 &0 \\ -5 & -2 &0 \\ 8& 3& 0 \end{bmatrix},\ det(A) =0\\ \\as\ one\ column\ is\ completely\ 0\\ \\so\ vectors\ are\ linearly\ dependent\\

so answer is option (c)

please upvote the answer.